### v2

# Cell as Point: One-Stage Framework for Efficient Cell Tracking

Abstract
Conventional multi-stage cell tracking approaches rely heavily on detection or segmentation in each frame as a prerequisite, requiring substantial resources for high-quality segmentation masks and increasing the overall prediction time. To address these limitations, we propose CAP, a novel end-to-end one-stage framework that reimagines cell tracking by treating Cell as Point. Unlike traditional methods, CAP eliminates the need for explicit detection or segmentation, instead jointly tracking cells for sequences in one stage by leveraging the inherent correlations among their trajectories. This simplification reduces both labeling requirements and pipeline complexity. However, directly processing the entire sequence in one stage poses challenges related to data imbalance in capturing cell division events and long sequence inference. To solve these challenges, CAP introduces two key innovations: (1) adaptive event-guided (AEG) sampling, which prioritizes cell division events to mitigate the occurrence imbalance of cell events, and (2) the rolling-as-window (RAW) inference strategy, which ensures continuous and stable tracking of newly emerging cells over extended sequences. By removing the dependency on segmentation-based preprocessing while addressing the challenges of imbalanced occurrence of cell events and long-sequence tracking, CAP demonstrates promising cell tracking performance and is 10 to 55 times more efficient than existing methods. The code and model checkpoints will be available soon.
1 Introduction
Figure 1: Our proposed CAP is an end-to-end trained framework leveraging the idea of Cell as Point to facilitate tracking cells efficiently. As (a) illustrates, instead of previous work requiring segmentation (SEG) or detection (DET) as a prerequisite for final tracking (TRA), CAP is able to track all cells within sequence frames in one stage. (b) shows that CAP reduces the inference time ($2.9$s) by approximately $\mathbf{10}$ to $\mathbf{55}$ times compared to previous works while maintaining high tracking performance of $0.93$. (c) demonstrates the cell tracking result predicted by CAP.
Cells are the fundamental units of organs in the human body [23]. Cell events are diverse, continuously changing, and intricate, involving cell migration (translocation), mitosis (division), and apoptosis (death) [44]. Cell tracking is the task of keeping track of each cell’s location and reconstructing cell lineages (mother-daughter relations), which help develop preventive therapies [63] and drug discovery [22]. Traditional methods of manual tracking demand significant resources and expertise to achieve reliable results. This complexity creates a pressing need for efficient, automated cell tracking systems.
Indeed, several existing works and platforms [28, 25, 26] have demonstrated relatively effective performance in achieving automated cell tracking. The taxonomies of strategies typically involve segmentation followed by linking (SegLnk) [24, 54], segmentation and linking (Seg&Lnk) [19, 43, 49], or even detection prior to the segmentation and linking stages (DetSegLnk [61] or DetLnkSeg [44, 27][52]. Also, frameworks [31, 32] often incorporate isolated data-preprocessing stages to represent cell tracking data. These multi-stage cell tracking pipelines can limit the overall performance of the framework, constrained by the errors that occur at earlier stages and are accumulated and amplified at subsequent stages [45]. Additionally, a multi-stage framework involving both detection/segmentation and tracking not only requires the tracking ground-truth but also demands additional labels, i.e., high-quality segmentation masks, to train the model. It is highly labor-intensive and time-consuming to produce the data annotation, rendering such methods impractical for clinical applications [86].
To overcome the data challenge for existing multi-stage frameworks, we propose a novel one-stage cell tracking framework, Cell as Point (CAP), as demonstrated in Fig. 1 (a), that treats each cell as a point to jointly track numerous cell points within a sequence. The one-stage CAP framework not only benefits from the efficient prediction procedure through a straightforward pipeline, shown in Fig. 1 (b), but also from its low requirement for data annotation, solely relying on tracking ground-truth (coarse masks and cell acyclic graphs) to represent the position and status of each cell. Specifically, by proposing cell point trajectory and visibility to represent the cell tracking data within a one-stage framework comprehensively, CAP offers an elegant solution to alleviate the issues of the complicated data-preprocessing and the dilemma of representation of cell decay. Another key innovation of CAP is the introduction of adaptive event-guided (AEG) sampling method and rolling-as-window (RAW) inference approach. Specifically, AEG sampling is a temporal data sampling strategy for cell microscopy imaging sequence data to remedy the imbalance of cell division occurrences in the dataset; RAW is an inference approach to capture and track new cells within a long sequence in an efficient way. To summarize, our key contributions include:
  • We propose a novel CAP cell tracking framework that is highly efficient in terms of inference time and data annotation requirements, enabling one-stage tracking that bypasses detection/segmentation to investigate cellular behaviors.
  • We address prevalent challenges in the cell sequence data by proposing AEG sampling to resolve imbalanced cell division occurrence, representing complex cell events for training purposes through cell point trajectory and visibility concepts within a one-stage framework, and employing a RAW inference approach to capture and track new cells in long sequences.
  • Our one-stage approach, CAP, achieves promising performance compared to existing multi-stage methods while alleviating the need for detection or segmentation.
2 Related Work
2.1 Cell Tracking
Cell tracking is a crucial task in biological imaging that typically involves the identification and temporal association of cell instances across image frames. Traditionally, this process includes two main steps: cell segmentation (detection) [1, 79, 55, 77] and cell linking [18]. The segmentation step identifies individual cells in each frame, while the linking step establishes temporal associations between these cells across frames, sometimes capturing complex dynamics such as cell division and migration [49, 52]. Nowadays, instead of traditional approaches such as Nearest Neighbors linking [67], Bayesian filters [66, 16, 33, 81, 34, 73], Viterbi [51, 50], and graph-based matching [5, 72, 3, 48], the advent of deep learning has transformed the landscape of cell tracking, leveraging advanced neural architectures for segmentation and tracking tasks. Deep learning approaches often employ complex models, e.g., CNN, RNN, GNN, and transformer, to combine multiple stages like detection [12, 10, 11, 46, 88, 30, 56, 78, 31, 32, 38], segmentation [82, 59, 9, 2, 49, 68, 7, 8, 64, 58, 17, 13, 57, 14, 71], and tracking into the frameworks. Several contemporary methods [62, 60, 30] aim to integrate contextual cell data into tracking models. ConvGRU [60] is introduced by Payer et al., capturing local cellular and inter-frame features. However, this method requires extensive annotated training data for all cell regions, making it labor-intensive. Motion and Position Map (MPM) is capable of representing detection and association at once, which is proposed by Hayashida et al. [31], marking an advancement in simplifying the pipeline for cell tracking. However, it still requires multiple stages, first generating the MPM followed by subsequent detection and tracking procedures.
Thus, although existing multi-stage methods, which are based on individual cell detection followed by linking/association or evolving segmentation masks over time, show great cell tracking performance, they heavily rely on segmentation as a preliminary step before tracking. These methods often require high-quality data annotation for training and can be computationally intensive, especially when dealing with high-resolution [83]. Additionally, these methods often involve complex pipelines that may not support end-to-end or one-stage training, posing challenges for both training and inference phases in terms of efficiency and scalability [45, 86]. Thus, to tackle these issues, an end-to-end one-stage framework that omits the intermediate segmentation or detection stage is essential.
2.2 Point Tracking
In the field of natural images, recent end-to-end video point-tracking frameworks have achieved commendable performance. Notably, TAP-Vid [20] pioneered tracking any point across video frames, establishing a benchmark, and claiming a baseline method. However, this method does not support tracking occluded points, limiting its applicability in cell-tracking scenarios. Even though PIPs [29] and PointOdyssey [87] methods take measures to address the issue of point occlusion in long sequences by tracking selected points using a fixed sliding window and resuming tracking from the last visible frame, these methods struggle with point tracking independently, which may not capture the dependencies between multiple points. OmniMotion [76] accounts for estimated correspondences within a canonical space by optimizing a volumetric representation for video sequences. Although innovative, this method involves test-time optimization, making it computationally expensive and less viable for clinical applications. CoTracker [40] represents significant advancements in the field of point tracking. It addresses the inaccuracies led by independent tracking among points by incorporating incremental point coordinates updates based on 4D correlation volumes, first proposed by RAFT [70] constructed for all pairs of points from point/object features. CoTracker3 [39] is advanced by involving a semi-supervised training methodology to reduce reliance on synthetic datasets and utilizing pseudo-labeling techniques to generate training labels from unannotated real-world videos.
However, the trackers designed for general-purpose applications are ineffective for cell tracking because the unique properties of cell behaviors, i.e., cell mitosis and apoptosis, have not been considered in the model construction. Specifically, cells are not only relative with position and occlusion but also potentially associated with their ancestors due to the regulation of cellular life decay and division. Therefore, they cannot simultaneously represent or predict both cell positions and division states with their ancestors or children. Moreover, regarding the publicly available cell tracking data [54, 52, 74, 41, 4], the diversity of cell types results in significant variations in cell division rates and sequence length across different datasets categorized by specimens. Imbalanced cell event occurrence [42] in a sequence can lead to the omission of complete division events (from the death of the mother cell to the birth of the daughter cell) during random sampling with a low division rate; the inference for long sequences typically requires more time or computational resources to predict cell locations and lineages at once. Thus, existing methods potentially result in the under-learning of division cases in the model training process and a higher resource requirement in the long-sequence inference procedure. Thus, the uniqueness of this time-lapse cell sequential image data poses the most significant challenge in developing a one-stage cell tracking framework to effectively track cells jointly.
3 Cell as Point (CAP)
Figure 2: Cell Point Trajectory and Visibility. M, D1, and D2 denote mother cell, daughter cell #1, and daughter cell #2. (x,y) in $L_{t}$ represent the location coordinates, and 00 or $1$ in $V_{t}$ respectively represent non-existing and existing cells. Three valid trajectory and visibility combinations: (a) M has not divided; (b) M has divided, and only D1 occurs in the $t$-th frame; (c) M has divided, and both D1 and D2 occur in the $t$-th frame.
Cell tracking aims to identify the characteristics of cell migration and division, thereby fully automating the prediction of specific cell position and division status changes in each frame of a time-lapse microscopy video. In our framework, CAP, every single cell is a point, tracking all cell points throughout the duration of a frame sequence $S=(I_{t})_{t=1}^{T}$ containing $T$ grayscale cell microscopy images $I_{t}\in\mathbb{R}^{H\times W}$. Cell point trajectories and visibilities are proposed to be utilized in the CAP framework to represent and predict cell locations and lineages. The cell point trajectory is constituted by the $N$ cell centroid point locations $L^{i}_{t}=(x^{i}_{t},y^{i}_{t})\in\mathbb{R}^{2}$, $i=1,\dots,N$, $t=1,\dots,T$, where $t$ is the time of a frame sequence. The cell point visibility of the cell $V^{i}_{t}\in\{0,1\}$ indicates whether the $i$-th cell point still exists in the frame sequence at time $t$; in other words, if the cell is non-existing or born due to mitosis and apoptosis. Since any cell division that happened in the event would have been that of the parent cell into two daughter cells [36]Fig. 2 demonstrates the combinations of cell point trajectory and visibility, representing various states of cells, i.e., migration, mitosis, apoptosis, and cell lineages (with two potential daughter cells). This simple yet effective representation method will efficiently aid the CAP framework in achieving one-stage efficient cell tracking.
3.1 Adaptive Event-Guided (AEG) Sampling
Figure 3: Overview of CAP framework. The training sequence $T_{s}$ is sampled from the entire frame sequence using AEG strategy. For each $T_{s}$, the cell point trajectories and visibilities are iteratively refined, starting from their initialization. For iteration $1,\dots,M$, updates are made to cell point trajectory $\hat{L}$ and tracking feature $F$ with cell point visibility $\hat{V}$ computed at the final ($M$-th) update. This figure illustrates a single iteration of the optimization process.
To reduce operational costs and compromise overfitting [6], instead of seeing the entire training sequence in a single instance, the model learns a continuous sub-sequence of $T_{s}$ frames for each time, where $0<T_{s}<T$. However, the imbalanced occurrence of cell division in the entire sequence would cause the model to detect cell division occurring rarely. To facilitate the effective training of models to understand the biological behavior of the completed cell division event while mitigating data imbalance [84, 47], we propose the Adaptive Event-Guided (AEG) sampling technique to balance the occurrences of cell division in a sequence of frames the model sees by determining the initial frame of $T_{s}$ learning a complete event.
The AEG method leverages the intrinsic dynamics of cell division within the frame sequences to guide the sampling process, enhancing the representational diversity of the training dataset and ensuring the model’s robustness to variations in cell behavior. Technically, as data loading iterates through the whole dataset by detecting the complete cell division processes (from the death of the parent cell to the birth of the daughter cell), an anchor is added at each time cell division occurs $\mathbb{T}_{anchors}$. The probability of applying AEG $P_{\text{AEG}}$ is adaptively calculated by $P_{\text{AEG}}=\frac{N_{\text{div}}\times T_{\text{div}}}{T}$, where $N_{\text{div}}$ denotes the total occurrence of cell division in the entire $T$-frame sequence, and $T_{\text{div}}$ denotes the duration of the completed cell division. When the random decimal $p<P_{\text{AEG}}$ indicates the usage of AEG, our CAP framework incorporates AEG sampling for selecting $T_{s}$, as illustrated at the top of Fig. 3; Otherwise, the starting frame of $T_{s}$ is randomly located in the $(0,T-T_{s})$ range.
For each training sequence, we crop a fixed-length frame sequence in accordance with the starting frame from the video. Occasionally, the process of cell division, resulting in the formation of new cells, may span multiple frames. Thus, the leverage of the AEG approach allows us to capture complete division processes within sampled sequences in data pre-processing. This fixed-length cropping is critical to maintaining uniformity in the input data, facilitating more stable and reliable learning outcomes.
3.2 Cell Joint Tracking
By regarding each cell as a point and leveraging the two-dimensional feature across time and a set of tracked cell points, we implement an effective method for cell tracking via the correlation between cell point trajectories and visibilities in one stage. The training process of CAP as a transformer $\Theta$ neural network [75], where $\Theta:G\mapsto O$. In order to improve the estimated cell trajectories and visibilities along with $M$ iterations, the trajectories and visibilities of $i$ points (cells) in a $t$-frame sequence are encoded into a grid of input tokens $G^{i}_{t}$. $O^{i}_{t}$ is the corresponding output of $\Theta$ for representing the refined trajectories and visibilities. The following sections elaborate on the transformer formulation visualized in Fig. 3.
Feature Extraction.
To initialize the tracking feature $F$ for the part of input tokens $G$, we first generate image features (feature map), $\phi(I_{t})$, for each frame in the training sequence, from which the tracking features, $F$, are distilled according to cell point locations. Specifically, the feature extractor is a convolutional neural network (CNN) trained end-to-end in the framework for computational efficiency. Leveraging the CNN feature extractor, the feature maps with $dim$ dimensions are generated from $I_{t}$, $\phi(I_{t})\in\mathbb{R}^{dim\times\left(\left\lfloor\frac{H}{k}\right\rfloor\times\left\lfloor\frac{W}{k}\right\rfloor\right)}$, where $k=4$ for computational efficiency, and the features are downscaled by average pooling with strides $s$ for multi-scaled features. The tracking features $F^{i}_{t=1}$ for $i$ are captured by bilinear interpolation from the location of cell points on the initial frame $L_{t=1}$ on the feature map $\phi(I_{t=1})$. The tracking feature vector $F^{i}_{t}\in\mathbb{R}^{dim}$ is initialized for the transformer input by broadcasting the tracking features in $t=1$ frame for all $T_{s}$ frames, which is refined and updated with $M$ iterations.
Correlation Volume.
To achieve the objective of associating cell point trajectories jointly, CAP employs correlation volume $C^{i}_{t}\in\mathbb{R}^{S}$, where $S=4$, proposed in RAFT [70]. These correlation volumes $C^{i}_{t}$ are derived by assessing the relationship between the tracking features $F^{i}_{t}$ and the image features $\phi(I_{t})$ in proximity to the estimated cell point location $\hat{L}^{i}_{t}$ of the trajectory on the $t$-th frame. Technically, each $C^{i}_{t}$ is the average over the correlation response in a $2^{s}\times 2^{s}$ grid:
$[C_{t}^{i}]_{\delta}^{s}$ $=\frac{1}{2^{2s}}\sum_{\delta}^{2^{s}}\sum_{\delta}^{2^{s}}\langle F^{i}_{t},\phi_{s}(I_{t})_{2^{s}\hat{x}+\delta,2^{s}\hat{y}+\delta}\rangle$ (1)
$=\langle F^{i}_{t},\frac{1}{2^{2s}}(\sum_{\delta}^{2^{s}}\sum_{\delta}^{2^{s}}\phi_{s}(I_{t})_{2^{s}\hat{x}+\delta,2^{s}\hat{y}+\delta})\rangle,$
where $s=1,\dots,S$ are the feature scales, and $\delta$ is the offset around cell points, $(\hat{x},\hat{y})$ is the estimated location coordinate of the cell point $\hat{L}_{t}^{i}$. The value of $[C_{t}^{i}]_{\delta}^{s}$ is computed as the inner product between the feature vector $F^{i}_{t}$ and $\phi_{s}(I_{t})$ at $\hat{L}_{t}^{i}$ pooled with kernel size $2^{s}\times 2^{s}$.
Iterative Updates.
In the transformer-based interactive updates, the input tokens express the concatenation of features: location $\hat{L}^{i}_{t}-\hat{L}^{i}_{1}$, visibility $\hat{V}^{i}_{t}$, tracking features $F^{i}_{t}$, and correlation volume of cell trajectories $C^{i}_{t}$. As for the output tokens $O(\Delta\hat{L},\Delta F)$, contain updates for location and refined tracking features: $O^{i}_{t}=(\Delta\hat{L}^{3i}_{t},\Delta F^{i}_{t})$, CAP estimates $3$ locations for each cell, following the representation illustrated in Fig. 2: the cell’s location and the locations of two potential daughter cells. In this way, CAP reserves space for each cell’s progeny by representing $L$ and $V$ through a $3$-column format. $M$-times updates the output tokens by applying the transformer $\Theta$ iteratively for a sequence $T_{s}$ as shown in Fig. 3. Specifically, CAP utilizes transformer $\Theta$ with cross-trajectory/time attention to update both the cell point trajectory, $\hat{L}$, and the tracking feature, $F$:
$\hat{L}^{(m+1)}=\hat{L}^{(m)}+\Delta\hat{L}~{}\text{and}~{}F^{(m+1)}=F^{(m)}+\Delta F,$ (2)
where $m=1,\dots,M$, and $m=1$ denotes initialization. At the last iteration of the update $M$, cell point visibility $V^{3i}_{t}$ for locations estimated $\hat{L}^{3i}_{t}$ is computed by $\hat{V}^{(M)}=\sigma(WF^{(M)})$, where $\sigma$ represents the sigmoid activation function and $W$ is a learned matrix of weights. By investigating the $3$ visibilities $V^{3i}_{t}$, CAP tells if the cell division occurred on the cell $i$ in the $t$-th frame.
Loss Function.
We supervised the transformer $\Theta$ on the trajectory and visibility between predicted and ground truth for every semi-overlapping window. The trajectory loss function $\mathcal{L}_{tra}$ is location regression summed over the iteration updates, and the visibility loss function $\mathcal{L}_{vis}$ is the cross entropy of binary visibility options:
$\mathcal{L}_{tra}(\hat{L},L)=\sum_{j=1}^{J}\sum_{m=1}^{M}\gamma^{M-m}\|\hat{L}^{(m,j)}-L^{(j)}\|,$ (3)
$\mathcal{L}_{vis}(\hat{V},V)=\sum_{j=1}^{J}\operatorname{CE}(\hat{V}^{(M,j)},V^{(j)}),$ (4)
where $\mathcal{L}_{vis}$ is updated at the $M$-th iteration, and $J$ is the total number of $T_{s}$ whose trajectory starts in the middle of the previous $T_{s}$.
Input: $\Theta$, $S_{\text{infer}}\leftarrow(I_{t})^{T_{\text{infer}}}_{t=1}$, $Q\leftarrow(x^{i},y^{i})$, $l_{\text{win}}$
Output: $L_{t}$, $V_{t}$, where $t\leftarrow 1,\dots,T_{\text{infer}}$
Initialization: current process frame $t_{\text{cur}}\leftarrow 1$
while $t_{\text{cur}}<T_{\text{infer}}$ do
       Obtain current window frames $S_{\text{w}}\leftarrow(I_{t})^{t_{\text{cur}}+l_{\text{win}}}_{t=t_{\text{cur}}}$
       $L_{\text{w}}$, $V_{\text{w}}\leftarrow\Theta(S_{\text{w}})$
       for $i\leftarrow 0$ to $l_{\text{win}}$ do
             $t_{\text{cur}}\leftarrow t_{\text{cur}}+i$
             $Q\leftarrow L_{t_{\text{cur}}}$
             if find new cell(s) then
                   Concatenate $Q$ and $\hat{L}_{t_{\text{cur}}}^{\text{new}}$
                   break
             end if
       end for
end while
Algorithm 1 Algorithm of RAW Inference
3.3 Rolling-as-Window (RAW) Inference
Metrics PhC-C2DH-U373 Fluo-N2DH-GOWT1 Fluo-C2DL-Huh7
GNN EmbedTrack KIT-GE CAP GNN EmbedTrack KIT-GE CAP GNN EmbedTrack KIT-GE CAP
$\mathrm{ES}$downarrow 214 0 3 0 49 39 45 18 34 0 34 31
$\mathrm{EA}$downarrow 371 44 50 0 140 119 111 120 85 733 77 71
$\mathrm{ED}$downarrow 0 0 0 0 1 0 0 2 2 0 0 0
$\mathrm{ESM}$downarrow 0 7 1 0 13 0 1 2 835 5 0 1
$\mathrm{FP}$downarrow 135 211 197 119 5 21 6 75 36 1392 41 101
$\mathrm{FN}$downarrow 25 32 25 0 28 40 32 46 4 636 0 1
$\mathrm{TRA}$uparrow 0.771 0.931 0.939 0.985 0.958 0.961 0.966 0.960 0.889 0.187 0.970 0.966
$\mathrm{IT}$downarrow 19.7 123.6 22.0 2.6 67.2 171.6 39.4 7.1 27.1 125.0 25.7 3.2
Metrics HeLa PC-3 3T3 RAW264
GNN EmbedTrack CAP GNN EmbedTrack CAP GNN EmbedTrack CAP GNN EmbedTrack CAP
$\mathrm{ES}$downarrow 2 0 3 2 0 1 6 2 107 186 68 67
$\mathrm{EA}$downarrow 385 211 82 52 43 11 613 603 949 1451 1475 1098
$\mathrm{ED}$downarrow 1 1 1 3 3 0 16 14 2 89 74 12
$\mathrm{ESM}$downarrow 13 1 0 12 1 0 3 4 4 21 28 17
$\mathrm{FP}$downarrow 62 31 60 33 0 35 100 201 250 239 669 647
$\mathrm{FN}$downarrow 342 218 70 0 42 7 9 589 709 1341 1259 898
$\mathrm{TRA}$uparrow 0.776 0.849 0.926 0.868 0.912 0.952 0.857 0.853 0.832 0.801 0.803 0.864
$\mathrm{IT}$downarrow 23.3 41.5 1.3 18.3 48.7 1.1 21.7 74.6 2.2 28.5 59.6 2.8
Table 1: Comparison Experiments with Baselines. $\mathrm{ES}$, $\mathrm{EA}$, $\mathrm{ED}$, and $\mathrm{ESM}$ are abbreviations of $\mathrm{EdgeSplit}$, $\mathrm{EdgeAdd}$, $\mathrm{EdgeDelete}$, and $\mathrm{EdgeSemantic}$, respectively. Our framework CAP is TRA only, while previous frameworks GNN, EmbedTrack, and KIT-GE are SEG+TRA illustrated in Fig. 1 (a). Due to space limitations, KIT-GE results on DeepCell datasets are included in the supplementary. Average inference time/seconds ($\mathrm{IT}$) and performance ($\mathrm{TRA}$) are visualized in Fig. 1 (b).
In cell tracking, the inference stage is crucial for producing the final tracking mask and establishing cell lineage relationships by analyzing predicted cell locations and division events. Existing works [68, 64, 78, 49, 8] often customize inference and post-processing stages with different configurations for each dataset, adapting to variations in sequence length. This data-specific tailoring limits their generalizability and broader applicability to other datasets. In contrast, our proposed Rolling-As-Window (RAW) inference approach in CAP simplifies and unifies these processes while maintaining effectiveness across varying sequence lengths, enabling high-quality inference regardless of computational constraints. Specifically, a consistent and high-quality inference result is achieved through a rolling loop, which sequentially extracts yet-to-be-inferred clips from the full sequence, processing them within a defined window for inference.
In a certain sequence with $T_{\text{infer}}$ frames, if the inference window length $l_{\text{win}}$ supported by either computational resources or architecture is shorter than $T_{\text{infer}}$, the first rolling of window frames $S_{\text{w}}=(I_{t})^{l_{\text{win}}}_{t=1}$ is passed into the trained transformer $\Theta$; otherwise, the whole sequence will be inferred at once. The RAW inference approach is summarized in Alg. 1, which will be triggered while the former scenario happens. Specifically, a $l_{\text{win}}$-length subsequence starting from $t=1$ is passed into $\Theta$ for prediction. The output, cell point trajectories $L_{t}$ and visibilities $V_{t}$, where $t=1,\dots,l_{\text{win}}$, are processed frame-by-frame to update query points $Q$ based on the $L_{t}$ of identified cells and check for cell division or the emergence of new cells. Whenever a new cell appears, the predicted location of this new cell $\hat{L}_{t}^{\text{new}}$ is incorporated into the updated query points $Q$ for the subsequent rolling of analysis by the formula:
$Q_{\text{r+1}}=\text{concat}(Q_{\text{r}},\hat{L}_{t}^{\text{new}})=\begin{bmatrix}Q_{\text{r}}\\ \hat{L}_{t}^{\text{new}}\end{bmatrix},$ (5)
where $r$ is the times of rolling for window inference. Subsequently, RAW segments a new window sequence starting from the frame where the new cell appeared, preparing it for the next application of $\Theta$. By adjusting the $l_{\text{win}}$ based on the capacity of computational resources and $T_{\text{infer}}$ value, our CAP is capable of performing efficient and high-quality inference on sequences of any length.
4 Experiments and Results
We conducted experiments to demonstrate the high efficacy and efficiency of CAP, our end-to-end, one-stage framework for cell tracking. Despite requiring minimal inference time and eliminating the need for additional segmentation masks, CAP achieves promising performance. Additionally, we perform comprehensive ablation studies to assess the effectiveness of individual CAP components.
4.1 Experimental Setup
Figure 4: Cell Trajectories. Each point represents a cell with distinct colors, indicating individual cell trajectories across time points.
Figure 5: Cell Divisions. (a) and (b) illustrate cell division events accurately detected by the CAP framework in the HeLa and 3T3 datasets, respectively.
Datasets.
We evaluate our proposed framework on DynamicNuclearNet Tracking in DeepCell [54] public database, HeLa, PC-3, 3T3, and RAW264 datasets divided by specimen, and ISBI cell tracking challenge [52, 74] public datasets, PhC-C2DH-U373 (U373), Fluo-N2DH-GOWT1 (GOWT1), and Fluo-C2DL-Huh7 (Huh7). These temporal data used for cell tracking, shot by different devices in various durations on diverse specimens, exhibit diverse morphologies and sequence characteristics. DeepCell contains $96$ sequences in total: $45$ for HeLa, $7$ for PC-3, $10$ for 3T3, and $34$ for RAW264, with $42$, $50$, $71$, and $45$ frames in each sequence, respectively. For the ISBI dataset, the U373 comprises glioblastoma-astrocytoma U373 cells in a $115$-frame sequence of size $520\times 696$. Dataset GOWT1 documents GFP-GOWT1 mouse stem cells into $92$ frames with a size of $1024\times 1024$ per sequence. Huh7 is a dataset of human hepatocarcinoma-derived cells in a sequence of $30$ frames with a $1024\times 1024$ size.
Implementation Details.
For the training process, we utilize $85$ training sequences from the DeepCell dataset and sequences $02$ from the ISBI dataset, using only tracking (TRA) ground truth (GT). During each iteration, we crop sequences of $T_{s}=24$ frames with a sliding window length of $l_{\text{slide}}=8$. We use pre-trained checkpoints from CoTracker [40], which was trained solely on synthetic RGB data. Since our input images are grayscale, we replicate the single channel three times to match the required RGB format. We use an end-to-end trained $2$-dimensional CNN architecture as the encoder to extract tracking features $F$. Specifically, it integrates bilinear interpolation and utilizes it to extract $S=4$ different scaled image features with $128$-channel output features concatenated to optimized dynamic feature extraction (Sec. 3.2). While training the transformer, we follow the pre-trained checkpoint that downsamples the resolution of frames to $384\times 512$ and choose $S=4$, $\Delta=3$ to calculate 4D multi-scaled correlation volumes (Sec. 3.2). More details can be found in the supplementary.
As for the testing (evaluation) process, we preprocess the first frame for each sequence using [64] to select query points $Q$ and start the process of window rolling. The evaluation sequence $S_{\text{infer}}$ follows the native resolution ($384\times 512$), which the dataset is trained on. We set $l_{\text{win}}=100$, which is mentioned in Sec. 3.3. We use the $\mathrm{TRA}$ evaluation criteria provided by ISBI [52, 74] and CTMC [4], which lay in the range $[0,1]$ that a higher score corresponds with a better performance and is calculated by $\mathrm{TRA}=1-\frac{\mathrm{min}(AOGM,AOGM_{0})}{AOGM_{0}}$, where $AOGM$ is the measure of acyclic oriented graphs matching [53], penalizing the number of transformations needed to transform the predicted tracking graph into the ground-truth tracking graph, e.g., merged cells ($\mathrm{EdgeSplit}$), missing links ($\mathrm{EdgeAdd}$), wrong links ($\mathrm{EdgeDelete}$), links with wrong semantics ($\mathrm{EdgeSemantic}$), false positives ($\mathrm{FP}$), and false negatives ($\mathrm{FN}$), and $AOGM_{0}$ is the $AOGM$ value for empty tracking results.
Baselines.
We compare our approach CAP with three methods, cell-tracker-gnn (GNN) [8], EmbedTrack [49], and KIT-GE [64], which have publicly available code and utilize the evaluation system provided by ISBI [52, 74] and acyclic oriented graphs matching (AOGM) measure [53]. For the ISBI dataset, to control for variables and ensure fairness, we train these methods on sequence $02$ and evaluate on $01$, following [64]. For the DeepCell data, we follow the experiments [54] and report the result. We report the measures involved in the cell tracking, $\mathrm{EdgeSplit}$, $\mathrm{EdgeAdd}$, $\mathrm{EdgeDelete}$, $\mathrm{EdgeSemantic}$, $\mathrm{FP}$, $\mathrm{FN}$, $\mathrm{TRA}$, and average inference time ($\mathrm{IT}$).
4.2 Results and Analysis
Tab. 1 presents the quantitative performance comparison of our method, CAP, and baseline methods, GNN, EmbedTrack, and KIT-GE on the ISBI dataset (U373, GOWT1, and Huh7) and DeepCell dataset (HeLa, PC-3, 3T3, and RAW264). Since $\mathrm{TRA}$ is derived from metrics $\mathrm{EdgeSplit}$, $\mathrm{EdgeAdd}$, $\mathrm{EdgeDelete}$, $\mathrm{EdgeSemantic}$, $\mathrm{FP}$, and $\mathrm{FN}$, we primarily focus on the $\mathrm{TRA}$ score for the model performance and $\mathrm{IT}$ for efficiency (highlighted in gray).
U373 dataset with a $115$-frame sequence not only does CAP outperform all GNN, KIT-GE, and EmbedTrack significantly, achieving a perfect score of 00 for $\mathrm{EdgeSplit}$, $\mathrm{EdgeAdd}$, $\mathrm{EdgeDelete}$, $\mathrm{EdgeSemantic}$, and $\mathrm{FN}$ with the highest $\mathrm{TRA}$ score of $0.985$, but also it has the smallest inference time of $2.6$s. In contrast, GNN, EmbedTrack, and KIT-GE exhibit higher error counts and inference time. The $\mathrm{TRA}$ of datasets HeLa, PC-3, and RAW264, $0.926$, $0.952$, and $0.864$, show that The CAP not only significantly leads other methods in terms of the $\mathrm{TRA}$ values but also utilizes markedly less inference time compared to other approaches. CAP still demonstrates comparable performance with the $\mathrm{TRA}$ score of $0.960$, $0.966$, and $0.832$ across previous works in datasets GOWT1, Huh7, and 3T3. The trade-off remains advantageous given that CAP directly eliminates the stages of segmentation or detection, thereby significantly reducing the volume of data labels required and inference time ($7.1$s, $3.2$s, and $2.2$s) by approximately $5.5$ to $39$ times, while it has a slightly lower $\mathrm{TRA}$ measure than the existing method. Fig. 4 highlights CAP’s robust cell trajectory tracking in time-lapse microscopy, while Fig. 5 demonstrates its strong capability in detecting cell divisions, accurately tracking mother and daughter cells across frames.
As for the extreme instability of EmbedTrack method on Huh7 with $\mathrm{TRA}=0.187$, our analysis suggests that the observed results stem from the method’s excessive reliance on the segmentation (SEG) stage and the usage of refined masks during the training of the segmentation model, while Huh7 contains only SEG golden truth (GT) masks without the more refined SEG silver truth (ST). The supplementary Appendix B illustrates more details about the difference between tracking GT mask, SEG GT, and SEG ST. This further substantiates an end-to-end one-stage cell tracking framework, which bypasses the segmentation or detection stage, enhances model stability, and raises the overall performance limits achievable by the model.
Overall, the CAP method demonstrates robust performance and constant high efficiency on inference and label usage across all datasets, consistently achieving high $\mathrm{TRA}$ scores and the lowest $\mathrm{IT}$s. Thus, CAP is a highly effective, efficient, and reliable method for cell tracking in a practical biomedical context.
4.3 Ablation Study
     AEG Sampling      $\mathrm{TRA}$ Score
     U373      GOWT1      Huh7
     ✗      0.556      0.941      0.942
     ✓      0.985      0.960      0.966
Table 2: Effectiveness of AEG Sampling. AEG sampling is built to balance the cell event (division) in a training sequence.
Effectiveness of AEG Sampling.
To mitigate the data imbalance in cell division in samples, AEG sampling Sec. 3.1 is an effective method we proposed to calibrate the sample experience in a complete cell division process. Tab. 2 illustrates that switching off AEG sampling decreases performance from $0.02$ to $0.43$.
Attention $\mathrm{TRA}$ Score
time joint U373 GOWT1 Huh7
0.974 0.903 0.944
0.985 0.960 0.966
Table 3: Effectiveness of Cell Joint Tracking. We compare CAP performance using time and cross-trajectory attention.
Effectiveness of Cell Joint Tracking.
By replacing the cross-trajectory attention layers with only time attention layers to keep the model size, Tab. 3 shows model performance declined by $0.015$ to $0.055$ across datasets. Thus, cell joint tracking plays an important role in CAP.
      Stride       $\mathrm{TRA}$ Score
      U373       GOWT1       Huh7
      8       0.531       0.609       0.853
      4       0.985       0.960       0.966
Table 4: Ablation on Feature Stride. CAP predicts more accurately with higher resolution features.
Ablation on Feature Stride.
Tab. 4 shows that the higher the resolution of the feature scaled, the more accurately CAP predicted. Therefore, we choose $s=4$ for scaling features during the training procedure (Sec. 3.2).
Additional ablations on DeepCell datasets are included in the supplementary.
5 Conclusion and Future Work
In this paper, we introduce the one-stage CAP (Cell as Point) framework, which eliminates the need for a separate segmentation or detection stage—a significant innovation in cell tracking. This approach substantially improves efficiency by reducing inference time by a factor of $10$ to $55$ and minimizing the need for high-quality data annotation. CAP effectively captures complex cell states and tracks cells in a single step, all while maintaining high efficiency. Additionally, it addresses common challenges in cell sequence data, such as cell event imbalance and long sequence inference. In conclusion, CAP not only offers considerable practical value in real-world applications due to its efficiency but also pioneers a novel approach in the field of cell tracking.
Limitations and Future Work.
Different datasets are captured by different devices, potentially resulting in variations in the luminance of cells presented in the time-lapse microscopy images; the datasets we utilized to evaluate our framework are annotated by various experts, so the quality of labels and annotation practices may vary; and different cell densities for datasets may cause disparities in performance. Consequently, a strategy to train a versatile model via our one-stage framework that could mitigate the gap and generalize to all datasets deserves future exploration.
Supplementary Material
This supplementary material complements the main manuscript by providing detailed information and additional support. It is structured as follows:
Appendix A: Additional Related Work
- Provides more related work about computing visual similarity and iterative updates proposed by RAFT.
Appendix B: Dataset Details
- Provides more detailed information and analysis on the ISBI and DeepCell datasets we utilized.
Appendix C: Experiment Details
- Offers the training and optimizer configuration details and evaluation metrics calculation details.
Appendix D: Additional Baseline Results
- Provides KIT-GE baseline method results on the DeepCell dataset as complementary for Tab. 1 in Sec. 4.2.
Appendix E: Additional Ablations
- Shows the effectiveness of adaptive event-guided sampling, cell joint tracking structure, and different feature strides in our proposed CAP framework.
Appendix A Additional Related Work
Optical flow refers to the pattern of apparent motion of objects, surfaces, or edges in a visual scene caused by the relative movement between an observer and the scene. In the natural imaging field, the optical flow algorithms, methods for estimating per-pixel motion in images [85], were studied through the lens of the color constancy equation [15, 37]. However, with the advancement in the deep learning field, more modern approaches [69], starting from FlowNet [21, 35] and DCFlow [80], introduced a 4D cost volume computation method and have shifted towards leveraging deep learning for high-accuracy optical flow estimation.
RAFT [70] is introduced by Teed and Deng, which further refined the DCFlow approach by incorporating incremental flow updates based on 4D correlation volumes constructed for computing the similarity between all pairs of pixels across two images [29], enabling the model to handle large displacements and complex motion patterns effectively. The model iteratively updates the flow field using a recurrent unit that performs lookups on the correlation volumes, leading to precise motion estimation [65].
A.1 Computing Visual Similarity
A.1.1 Correlation Volume
Given two feature maps $F_{1}\in\mathbb{R}^{H\times W\times dim}$ and $F_{2}\in\mathbb{R}^{H\times W\times dim}$ with output $dim$-dimensions, which are features extracted from two input images $I_{1}$ and $I_{2}$, by feature extractor $g_{\theta}$:
$F_{1}=g_{\theta}(I_{1})\text{ and }F_{2}=g_{\theta}(I_{2}).$ (6)
RAFT computes a 4D correlation volume $C$ that captures the similarity between all pairs of pixels across the two images. The correlation between a pixel at position $(i,j)$ in $F_{1}$ and a pixel at position $(k,l)$ in $F_{2}$ is calculated as the dot product of their feature vectors:
$C_{ijkl}$ $=\langle F_{1}(i,j),F_{2}(k,l)\rangle$ (7)
$=\sum_{h}F_{1}(i,j,h)\cdot F_{2}(k,l,h).$
The correlation volume is a 4D tensor of dimensions $H\times W\times H\times W$, where $H$ and $W$ are respectively the height and width of feature maps $F$s.
A.1.2 Correlation Pyramid
RAFT constructs a correlation pyramid by applying average pooling to the last two dimensions of the 4D correlation volume to capture multi-scale feature information. Specifically, this pooling is performed with kernel sizes of $1$ $2$, $4$, and $8$, and a set of multi-scale correlation volumes $C^{k}=\{C^{1},C^{2},C^{3},C^{4}\}$ is generated. Thus, each volume $C^{k}$ in the pyramid has dimensions $H\times W\times\frac{H}{2^{k}}\times\frac{W}{2^{k}}$.
A.2 Iterative Updates
The update operator estimates a sequence of flow estimations $\{f_{1},\dots,f_{n}\}$ with initializing the start point $f_{0}=0$. For each iteration, it produces $\Delta f$ as the direction that should be applied to the current estimation:
$f_{k+1}=\Delta f+f_{k+1}.$ (8)
Specifically, given the current flow estimation $f_{k}$, the correlation volume pyramid is retrieved and undergoes processing through two convolutional layers to extract relevant information. The updated information is comprised of correlation, flow, and context features and is updated over time as a feature map concatenation.
Appendix B Dataset Details
B.1 ISBI Dataset
Figure 6: Comparison of Quality and Requirement of Data Annotation. Fluo-C2DL-Huh7 does NOT have segmentation ST. Instead of previous works utilizing all types of mask information with original images in the figure, CAP only utilizes the Tracking GT masks with original images (in the orange block).
Figure 7: Visualization of the DeepCell Dataset. HeLa, PC-3, 3T3, and RAW264 have different cell densities, sizes, and luminances.
To support the claim in Sec. 4.2, Fig. 6 illustrates datasets we leveraged, PhC-C2DH-U373, Fluo-N2DH-GOWT1, and Fluo-C2DL-Huh7 [52, 74], with the difference between tracking (TRA) GT mask, segmentation (SEG) GT, and segmentation (SEG) ST if available. No segmentation ST is available specific to dataset Fluo-C2DL-Huh7. We observe that SEG GT and ST are much more refined and high-quality than TRA GT masks, requiring more resources and manpower. Fig. 6 also demonstrates the comparison of data annotation requirements between the previous works and our framework CAP that, instead of previous works leveraging all TRA GT masks, segmentation GT, and ST if available, CAP utilizes the TRA GT masks. Thus, in addition to being a highly efficient framework for inference time, the CAP also excels in efficiency with respect to both the quality and quantity of labels required, which is attributed to the one-stage characteristic of the framework.
Furthermore, detailed information on data collection and the characteristics of each dataset are described below:
  • The PhC-C2DH-U373 dataset comprises active glioblastoma-astrocytoma U373 cells on a polyacrylamide substrate. Each sequence is captured by a Nikon microscope with an object lens Plan Fluor DLL 20x/0.5 for a relatively long duration of 15 minutes, totaling $115$ frames with the size of $520\times 696$. The cells are relatively sparse in frames. The distinct feature of this dataset, compared to others, is its higher grayscale background.
  • Dataset Fluo-N2DH-GOWT1 documents the activity of GFP-GOWT1 mouse stem cells over a 5-minute period as observed by Leica TCS SP5 microscope with objective lens Plan-Apochromat 63x/1.4 (oil), segmented into $92$ frames with a size of $1024\times 1024$ per sequence. The target cells for tracking are notably active in cell events and exhibit a relatively high density within the frames.
  • Fluo-C2DL-Huh7 is a dataset of human hepatocarcinoma-derived cells expressing the fusion protein YFP-TIA-1. The Nikon Eclipse Ti2 microscope with a lens CFI Plan Apo Lambda 20x/0.75 is the shooting device to capture the cell migration, death, and division in a duration of 15 minutes with $30$ frames for each sequence. Cells in this dataset exhibit a relatively low frequency of cell events; However, due to the smaller number of frames recorded within the same time period, some cells exhibit significantly larger movements than others within a sequence. Cells are relatively small and have high density in frames.
B.2 DeepCell Dataset
The DynamicNuclearNet Tracking in the DeepCell [54] database consists of HeLa, PC-3, 3T3, and RAW264 datasets, which respectively contains cell type HeLa-S3, PC-3, NIH-3T3, and RAW 264.7, visualized in Fig. 7. These four cell types are mammalian cell lines that were acquired from the American Type Culture Collection [54]. The cells are cultured in Dulbecco’s modified Eagle’s medium (RAW 264.7 and NIH-3T3) or F-12K medium (Hela-S3 and PC-3) supplemented with the necessary drugs, e.g., L-glutamine, penicillin, for all other cells.
There are $5$, $2$, $1$, and $3$ evaluation sequences for datasets HeLa, PC-3, 3T3, and RAW264, respectively. In the main manuscript, we report the average metric values for each dataset. In the following additional ablation study Appendix E in this supplementary, we report the detailed metric values for each sequence.
The detailed information on each cell type is illustrated below:
  • HeLa includes $77180$ objects, $2234$ tracks, and $189$ cell divisions, with a relatively high cell density. Each sequence is segmented into $42$ frames, lasting $6$ minutes, containing cell behaviors, i.e., cell migration and division. For each frame in the sequences, they are $540\times 540$ with a pixel size $0.65$.
  • PC-3 is the smallest subset in DynamicNuclearNet, with $5051$ objects, $159$ tracks, and only $5$ divisions, captured in a $5$-minute duration. All sequences are segmented into $50$ frames with a size $584\times 584$ and pixel size $0.55$.
  • 3T3 cells are fibroblasts from mouse embryonic tissue. In this paper, this cell line contributes $62983$ objects, $1413$ tracks, and $250$ divisions, offering insights into cell growth and cycle dynamics. There are $71$ frames in each sequence, and each frame is $512\times 512$ with a pixel size of $0.65$.
  • The RAW264 cell line, derived from murine macrophages, includes $335049$ objects, $8642$ tracks, and $1322$ divisions, making it the largest subset in our study. Each sequence has $45$ frames with a size of $540\times 540$ with a pixel size of $0.65$. The cell density in this dataset is the highest, with the smallest cell size.
Hyperparameter Value
Learning rate scheduler OneCycleLR
Upper learning rate 5e-4
Weight decay 1e-5
Epsilon 1e-8
Optimizer AdamW
Percentage of the cycle increasing the learning rate 0.05
Cycle momentum False
Annealing strategy linear
Table 5: Hyperparameter Configuration for CAP.
Metrics HeLa PC-3 3T3 RAW264
KIT-GE CAP KIT-GE CAP KIT-GE CAP KIT-GE CAP
$\mathrm{ES}$downarrow 6 3 29 1 128 107 37 67
$\mathrm{EA}$downarrow 57 82 56 11 783 949 585 1098
$\mathrm{ED}$downarrow 1 1 0 0 14 2 106 12
$\mathrm{ESM}$downarrow 17 0 0 0 10 4 24 17
$\mathrm{FP}$downarrow 70 60 13 35 256 250 718 647
$\mathrm{FN}$downarrow 67 70 1 7 698 709 603 898
$\mathrm{TRA}$uparrow 0.923 0.926 0.937 0.952 0.820 0.832 0.876 0.864
$\mathrm{IT}$downarrow 14.4 1.3 9.1 1.1 19.1 2.2 96.7 2.8
Table 6: Comparison Experiment with Baseline. Here is a complementary for Tab. 1 in Sec. 4.2.
Appendix C Experiment Details
C.1 Training and Optimizer
By utilizing the pre-trained CoTracker [40] checkpoint, we freeze the CNN feature extractor $\phi$ and use the AdamW optimizer to train the proposed CAP model. We present hyperparameter configuration for the optimizer in Tab. 5.
C.2 Evaluation Metrics
We follow the evaluation metrics of the operations to correct the situations of merged cells, missing links, wrong links, links with wrong semantics, false positives, and false negatives: $\mathrm{EdgeSplit}$ ($\mathrm{ES}$), $\mathrm{EdgeAdd}$ ($\mathrm{EA}$), $\mathrm{EdgeDelete}$ ($\mathrm{ED}$), $\mathrm{EdgeSemantic}$ ($\mathrm{ESM}$), $\mathrm{FP}$, and $\mathrm{FN}$. $\mathrm{TRA}$ measure [52, 74] is calculated from AOGM:
$\mathrm{TRA}=1-\frac{\mathrm{min}(AOGM,AOGM_{0})}{AOGM_{0}},$ (9)
$AOGM=w_{\mathrm{ES}}\mathrm{ES}+w_{\mathrm{EA}}\mathrm{EA}+w_{\mathrm{ED}}\mathrm{ED}+$
$w_{\mathrm{ESM}}\mathrm{ESM}+w_{\mathrm{FP}}\mathrm{FP}+w_{\mathrm{FN}}\mathrm{FN}.$ (10)
AOGM is the weighted sum of the executed operations considered as the cost of transforming ($\mathrm{ES}$, $\mathrm{EA}$, $\mathrm{ED}$, $\mathrm{ESM}$, $\mathrm{FP}$, and $\mathrm{FN}$) the computed graph into the reference one. and $\mathrm{TRA}$ used in the ISBI cell tracking challenge. The details of $\mathrm{ES}$, $\mathrm{EA}$, $\mathrm{ED}$, $\mathrm{ESM}$, $\mathrm{FP}$, and $\mathrm{FN}$ are elaborated in the following:
AEG Sampling HeLa PC-3 3T3 RAW264
01 02 03 04 05 01 02 01 01 02 03
0.898 0.908 0.883 0.921 0.878 0.952 0.924 0.829 0.804 0.828 0.843
0.962 0.958 0.921 0.900 0.891 0.980 0.924 0.832 0.824 0.868 0.900
Table 7: Effectiveness of AEG Sampling. AEG sampling is built to effectively balance the cell event (division) in a training sequence.
Attention HeLa PC-3 3T3 RAW264
time joint 01 02 03 04 05 01 02 01 01 02 03
0.853 0.949 0.861 0.921 0.906 0.925 0.947 0.812 0.805 0.845 0.821
0.962 0.958 0.921 0.900 0.891 0.980 0.924 0.832 0.824 0.868 0.900
Table 8: Effectiveness of Cell Joint Tracking. We compare CAP performance using time and cross-trajectory attention.
Stride HeLa PC-3 3T3 RAW264
01 02 03 04 05 01 02 01 01 02 03
8 0.884 0.919 0.806 0.936 0.853 0.105 0.274 0.387 0.610 0.622 0.758
4 0.962 0.958 0.921 0.900 0.891 0.980 0.924 0.832 0.824 0.868 0.900
Table 9: Ablation on Feature Stride. CAP predicts more accurately with higher resolution features.
$\mathrm{\mathbf{ES}}$ is the operation for non-split vertices that the computed markers with more than one reference marker assigned.
$\mathrm{\mathbf{EA}}$ is the operation for missing links that the cell lineage relations are not indicated.
$\mathrm{\mathbf{ED}}$ is the operation for redundant links that the fake cell lineage relations are indicated.
$\mathrm{\mathbf{ESM}}$ is the operation for altering the edge semantics, that is, the difference in semantics between the reference graph and the induced subgraph.
$\mathrm{\mathbf{FP}}$ is the situation in the extra detected objects (i.e., computed markers without any reference marker assigned).
$\mathrm{\mathbf{FN}}$ is the situation in the missed objects (i.e., reference markers not assigned to any computed marker).
Appendix D Additional Baseline Results
Tab. 6 provides additional baseline results as a complementary analysis to Tab. 1 in Sec. 4.2., further demonstrating the effectiveness of our CAP framework. It shows the performance comparison of our proposed CAP framework against KIT-GE across four datasets, HeLa, PC-3, 3T3, and RAW264, in the DeepCell. CAP consistently achieves higher TRA (tracking accuracy) scores than KIT-GE in HeLa, PC-3, and 3T3 datasets, demonstrating its superior tracking capability. Furthermore, CAP significantly reduces IT (inference time) across all datasets, being 10×10$\times$10 × to $\times8$ faster than KIT-GE, highlighting its efficiency. Overall, CAP delivers superior tracking accuracy with significantly lower inference time, reinforcing its effectiveness as a one-stage tracking solution compared to the multi-stage KIT-GE framework.
Appendix E Additional Ablations
Following the initial assessment of adaptive event-guided (AEG) sampling, cell joint tracking structure, and feature stride on the ISBI dataset in Sec. 4.3, we extended our ablation studies to the DeepCell dataset to validate the generalizability and effectiveness of these components in our CAP framework. The findings, presented in Tab. 7Tab. 8, and Tab. 9, underscore the significant impact of each component on the framework’s performance across various cell types and conditions.
E.1 Effectiveness of AEG Sampling
Tab. 7 shows that adaptive AEG sampling effectively balances cell event occurrences during training sequences, which enables AEG sampling to generally improve performance across all cell types. For instance, in the HeLa dataset, AEG sampling boosts the scores from $0.898$ to $0.962$ in the best case, indicating a substantial enhancement in the model’s ability to handle imbalanced data effectively. This improvement is consistent across other cell types like PC-3 and RAW264, affirming AEG sampling’s role in optimizing event-centric learning.
E.2 Effectiveness of Cell Joint Tracking
As represented in the main manuscript, we replace the cross-trajectory attention layers with only time attention layers to keep the model size. As illustrated in Tab. 8, the integration of both time and joint attention mechanisms in cell joint tracking significantly enhances tracking accuracy. The dual attention strategy not only improves temporal coherence but also reinforces the relational context between trajectories, evident from the increased scores in datasets such as HeLa, where the combined attention boosts the performance from $0.853$ to $0.962$ in the best cases. This enhancement highlights the importance of sophisticated tracking mechanisms in complex cellular environments.
E.3 Ablation on Feature Stride
The impact of feature stride on model accuracy is depicted in Tab. 9. Reducing the stride from $8$ to $4$ leads to a marked improvement in the resolution of feature maps, which in turn enhances the model’s predictive accuracy. For example, in the HeLa dataset, reducing the stride results in an increase from $0.884$ to $0.962$ in the best scenarios. This finding suggests that higher resolution features are critical for accurate cell tracking, particularly in datasets with intricate cell dynamics and small cell sizes like 3T3 and RAW264.